OpenStudy (anonymous):

A train starting from rest is moving with a uniform acceleration of 0.2 m/s^2 for 5 mins , Calculate the speed acquired and the distance traveled in this time... @Shane_B @angela210793 @lgbasallote @baddinlol

OpenStudy (ganpat):

in what time ?? i mean how much time.. ??

OpenStudy (ganpat):

a= (v - u)/ t.. this might help..

OpenStudy (anonymous):

\[\huge NOW??\]

OpenStudy (anonymous):

you can solve it further @Ganpat

OpenStudy (ganpat):

k, so t = 5 min.. a= 0.2, u = 0 and v = ?? so, 0.2 = (v-0)/ (5* 30) v = 150 * 0.2 V = 75 m /s..

OpenStudy (anonymous):

okay is that alllllll

OpenStudy (anonymous):

@baddinlol is he correct

OpenStudy (anonymous):

@Ganpat, doesn't time have to be in seconds?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

You used it in minutes..

OpenStudy (anonymous):

Your answer

OpenStudy (anonymous):

Wait let me solve it.

OpenStudy (anonymous):

For speed required

OpenStudy (ganpat):

@baddinlol: yes, so i multiplied t by 60..

OpenStudy (anonymous):

Use v = u + at

OpenStudy (ganpat):

i have done calculation mistake.. sorry

OpenStudy (anonymous):

ohh...

OpenStudy (ganpat):

k, so t = 5 min.. a= 0.2, u = 0 and v = ?? so, 0.2 = (v-0)/ (5* 60) v = 300 * 0.2 V = 150 m /s.. this is right

OpenStudy (anonymous):

ooooooooppppppssssssssss

OpenStudy (anonymous):